A lawyer charges $120for every 3/4hours. What is the lawyers rate per hour

Answer by stanbon(75887)

A lawyer charges $120for every 3/4hours. What is the lawyers rate per hour
-------
Use a proportion::
rate/time = rate/time
----
r/1 = 120/(3/4)
r = 40*4 = $160 per hour
-----------------------------
Cheers,
Stan H.
-------------------